Paul Garnett has an investment opportunity that promises cash payments of $25,000 at the end of each year for eight years. How much should he be willing to invest in the opportunity if he wants to earn 10.2% APR compounded quarterly?
Answer #1

EAR rate due to quarterly compounding = (1+10.2%/4)4 -1 = 10.5968%
Amount to be invested using annuity formula = PMT*(1-(1+r)-n)/r = 25000*(1-(1+10.5968%)-8)/10.5968% = 130,524.12

IF annual rate = 10.2%
Amount to be invested using annuity formula = PMT*(1-(1+r)-n)/r = 25000*(1-(1+10.2%)-8)/10.2% = 132,407.60

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well